# Efficient Response Generation Strategy Selection for Fine-Tuning Large Language Models Through Self-Aligned Perplexity

### v2

@@ Abstract
Fine-tuning large language models (LLMs) typically relies on producing large sets of input-output pairs. Yet for a given question, there can be many valid outputs. In practice, these outputs are often derived by distilling knowledge from teacher models, and they can vary depending on the specific teacher model or prompting strategy employed. Recent findings show that how these training outputs are generated can significantly affect the performance of the fine-tuned model, raising an important question: how do we pick the best data generation method from among numerous possibilities? Rather than exhaustively training and evaluating on each candidate, this paper proposes a scalable approximate method that assesses a small subset of generated data to estimate its suitability for a specific target LLM. Our central idea is that effective outputs should be familiar to the target LLM. While previous work measures familiarity with perplexity, we find that perplexity might be suboptimal in characterizing “familiarity” through theoretical analysis and practical observations. To address this, we introduce self-aligned perplexity, a novel metric capturing how closely candidate outputs adhere to the target LLM’s own style and reasoning patterns. In this way, we can identify the most effective generation strategy on a small sample, then apply it to produce the complete training set. We demonstrate that training on data generated by the chosen method yields significant improvements across diverse reasoning-focused benchmarks.
@@ 1 Introduction
Prior research has explored improving responses by adding details or rationales, such as structuring ground truth step by step (Hsieh et al., 2023; Ranaldi & Freitas, 2024), incorporating rationales, or enriching responses with additional information (Zhang et al., 2024; Kang et al., 2023; Li et al., 2022). However, recent studies (Ren et al., 2024; Yang et al., 2024) suggest that more details or converting responses to step by step style do not always improve performance and that alignment with the LLM’s linguistic style is crucial.
@@
From our experiment, we observe that, no single response generation strategy works universally across tasks. Thus, we need to creating a method to find out the most effective way to generate response for each task, rather than use a single method for all of the tasks.
@@
Some works (Xu et al., 2024; Kim et al., 2024) attempt to predict the effectiveness of response generation methods by evaluating the entire training dataset. They generate full training datasets using each method and then estimate training effectiveness based on scores computed via algorithms or reward models. However, these approaches are computationally expensive and not scalable.
@@
However, can we predict the effectiveness of each data generation methods efficiently? We observe an interesting phenomenon that each response generation method produces responses with a consistent style, meaning that a small subset of generated examples can effectively represent the entire dataset. Based on this assumption, we propose an efficient ranking pipeline that evaluates a limited number of samples (e.g., 10) to assess the performance of each response generation strategy. This approach uses an alignment estimation function to assign scores to each strategy, enabling us to identify the best-performing method without the need for a full-dataset evaluation.
@@
Previous research (Ren et al., 2024) used perplexity to measure a model’s familiarity with candidate question-answer pairs, proposing that lower-perplexity responses for the same input tend to yield better training performance. However, we found several cases where perplexity-based filtering was ineffective. For instance, responses structured in a step-by-step or redundant language style often have low perplexity but do not necessarily improve training outcomes.When examining the initial response from the target LLM, we note that on some tasks, the probability of the model producing a step-by-step or redundant response in its initial prediction is very low, even though these responses have low perplexity. These findings suggest that perplexity can be easily ”hacked” by response style. Consequently, traditional perplexity alone is insufficient for selecting the best response generation strategy.
@@
To address this, we propose self-aligned perplexity, a refined metric for measuring a model’s familiarity with target responses. The key idea is that a model is most familiar with the data it generates itself. Leveraging this, we modify perplexity computation by incorporating model-generated responses as in-context examples. Specifically, we first have the model produce an initial response, which is then appended to the question as an in-context example. A prompt enforce the model to pay attention to this example when computing perplexity, thereby altering the probability estimation of the candidate response. If the target response deviates significantly from the model’s own generated response—the one it is most familiar with—the model assigns it a lower probability, increasing its perplexity. Our experiments show that self-aligned perplexity outperforms traditional perplexity in selecting effective data generation strategies.
@@
In our experiments, we observe a strong correlation between the proposed indicator and the ranking of training dataset performance. Furthermore, we construct a pool of answer generation strategies and demonstrate that applying our selection criterion leads to significant performance gains compared to the baselines.
@@ 2 Related Works
There has been extensive research into what types of data yield the best training outcomes for large language models (LLMs). Previous studies have identified several factors that positively influence model training, such as adding complexity (Xu et al., 2023), adding details (Zhang et al., 2024; Kang et al., 2023; Li et al., 2022), adding diversity (Luo et al., 2023), augmenting ground-truth answers in a step-by-step manner (Hsieh et al., 2023; Ho et al., 2022; Magister et al., 2023; Fu et al., 2023; Ranaldi & Freitas, 2024), and ensuring correctness (Trinh et al., 2024; Ranaldi & Freitas, 2024). However, in practice, these metrics are challenging to measure for a given dataset, making it difficult to determine the quality of training data based on these criteria. Ren et al. (2024) found that familiarity, measured by perplexity, significantly impacts model training. They argue that language models are more familiar with content generated by other language models, which explains why synthetic data is often more effective for training compared to human-generated data for educating LLMs.
@@
Perplexity has been widely used for different purpose in prior research. Perplexity has been used to select prompts (Gonen et al., 2022), showing that prompts with lower perplexity generally lead to better performance in question-answering tasks. It has also been used for selecting pretraining datasets (De la Rosa et al., 2022), detecting AI-generated content (Xu & Sheng, 2024; Hu et al., 2020), and selecting instruction-tuning data from a database (Mekala et al., 2024). Li et al. (2024) modify the perplexity score and propose “IDF” (Instruction Following Difficulty), which is used to select a small pool of challenging data from the original dataset for efficient training. Researchers hypothesized that higher perplexity reflects more challenging data, which can be beneficial for teaching LLMs new knowledge. Unlike these studies, our focus is on finding the best strategy to generate the target responses (y) for a given x.
Two recent works have explored optimal strategies for generating responses, but their approaches diverge from ours. In (Xu et al., 2024), the authors introduce the Compatibility-Adjusted Reward (CAR) to rank training datasets; however, they limit their experiments to two instruction-following datasets and do not consider multiple meta-prompts. Similarly, (Kim et al., 2024) identifies the key factors influencing training outcomes and employs principal component analysis (PCA) to fit a combination of these factors in order to predict the training performance for the same datasets. In contrast, our work presents several key distinctions. First, unlike the previous work, we evaluate only a small subset of generated data rather than complete datasets, enabling more efficient analysis. Second, while previous studies rely on composite metrics that require hyperparameter tuning—potentially risking overfitting on limited data diversity—we employ a single, robust metric validated across diverse tasks. Third, instead of ranking teacher models using a single meta-prompt, we rank models paired with diverse meta-prompts. Lastly, rather than solely demonstrating ranking effectiveness, our method also enhances overall accuracy.
@@ 3.1 Problem Definition
Let be a set of candidate answer-generation strategies, where each strategy produces a response for an input . Our goal is to select the strategy that yields the most effective training data for fine-tuning a target model . Since generating the full dataset via the API for every strategy is costly, we evaluate a small subset of size () to estimate how well each strategy’s outputs align with .
@@ 3.2 The familiarity hypothesis
The work in Ren et al. (2024) suggests that if the model is more “familiar” with a given response, then the model can learn better with the given response. In Ren et al. (2024), perplexity, which is equivalent to the likelihood of generating a response with the model, is used to measure this familiarity score. In our study, we argue that perplexity might be sub-optimal to measure the familiarity. We suggest that the familiarity can be more precisely (at least theoretically) measured by the following equation:
where is a semantic similarity measure between and a sample response drawn from the model . In plain language, it quantifies how similar a candidate response is to the range of answers that the model might generate. It is straightforward to demonstrate that when , i.e., when only if is exactly identical to , the function becomes equivalent to the likelihood , and hence equivalent to perplexity. Theoretically, using perplexity as a surrogate to measure familiarity fails to account for the variety of responses that may be semantically equivalent to a candidate response, thereby underestimating the familiarity. In practice, this results in assigning an excessively high perplexity to a good candidate response that the model might actually be familiar with, as evidenced by our empirical study in section 6.1 and examples in Table 1.
| Model | Task | Prompt | Accuracy | Perplexity | Self-algined PPL |
| Mistral7B | GSM8K | Step by Step | 0.627 | 2.09 | 2.26 |
| Claud Answer Directly | 0.592 | 1.96 | 2.45 | ||
| Llama3 | MATH | Step by Step | 0.56 | 1.88 | 1.91 |
| Claud Answer Directly | 0.495 | 1.84 | 2.02 | ||
| Qwen2.5 | API-Bank | Step by Step | 0.41 | 3.02 | 2.98 |
| Claud Answer Directly | 0.461 | 3.14 | 2.90 |
@@ 3.3 Self-Aligned Perplexity
To address the issue described above, we introduce an alternative surrogate for measuring familiarity by leveraging the concept of in-context learning. Specifically, we use the model’s initial prediction, , as an in-context example of style and then evaluate the likelihood of a candidate response conditioned on this example in the prompt. This approach encourages the model to assign a higher probability to a response if its style closely aligns with that of the model’s generated response.
Traditional Perplexity.
For a candidate response of length , traditional perplexity is defined as:
Self-Aligned Perplexity.
To address this, we use ’s initial prediction, , as an in-context style example. For each input in a subset :
1.
Generate . 111 may generate incorrect responses, which can sometimes cause to mislead the in-context perplexity calculation. To mitigate this, we employ a filtering mechanism to remove incorrect . According to our experimental results, since is primarily used to guide style, its correctness is less critical. Hence, the correctness filter is optional. For more details, please refer to Appendix A.4 and A.7.
2.
For candidate , include (or when ) in the prompt to provide style guidance.
3.
Compute the perplexity of conditioned on this prompt.
Formally,
This self-aligned measure penalizes responses that deviate from ’s own style. For more details, please read Appendix A.1
Selection Criterion.
For each strategy , we calculate:
and select:
Table 1 shows that self-aligned perplexity more reliably selects the optimal strategy across three tasks (GSM8K, MATH, and API-Bank). For example, on GSM8K, Mistral7B achieves its best accuracy with the step-by-step strategy—correctly favored by self-aligned perplexity but not by standard perplexity.
@@ 4 Benchmark Construction
In the section, we shows how do we use different strategies (distinct prompts and teacher LLMs) in generating the high-quality responses with different styles.
@@ 4.1 Target LLMs and APIs
We use Mistral-7B-instruct-V2 (Jiang et al., 2023), Llama3-instruct (Dubey & Abhinav Jauhri, 2024) and Qwen-2.5-7B-Instruct(Qwen et al., 2025) as the target language models . In this paper, we refer to Llama3-instruct, Mistral-7B-instruct-V2, and Qwen-2.5-7B-Instruct as Mistral7B, Llama3, and Qwen2.5, respectively. We use GPT-4o, MiniGPT-4o, and Claude 3.5 APIs as teacher models for response generation. Specifically, we use gpt-4o-mini-2024-07-18 and gpt-4o-2024-08-06 (OpenAI, 2023) from OpenAI, and claude-3-5-sonnet-20240620 (Anthropic, 2023) from Anthropic.
@@ 4.2 Datasets
We use English reasoning datasets referenced in the technical reports of LLaMA3 (Dubey & Abhinav Jauhri, 2024), Mistral (Jiang et al., 2023), and Qwen-2.5 (Qwen et al., 2025) (the three target models in our experiments). We select datasets with at least 650 examples that can be evaluated via accuracy. If a dataset lacks sufficient training data, we reconstruct it to contain at least 400 training, 50 validation, and 200 testing examples.
For datasets with subcategories (e.g., MATH, MMLU, MMLU_PRO, API_BANK, AGIEVAL), we choose the most challenging subcategory (i.e., with the lowest reported accuracy). For example, we include moral scenarios from MMLU, Professional Law from MMLU_PRO, Level 3 problems from API_BANK, geometry from MATH, and LogicQA from AGIEVAL; we also incorporate the Algebra subcategory from MATH as in Ren et al. (2024).
@@
In total, our datasets include: Mathematics: GSM8K (Cobbe et al., 2021), MATH (Algebra) and MATH (Geometry) (Hendrycks et al., 2021); Commonsense reasoning: PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), Hellaswag (Zellers et al., 2019), and ECQA (Aggarwal et al., 2021); Reading comprehension: BoolQ (Clark et al., 2019) and SQuAD (Rajpurkar et al., 2016); Aggregated benchmarks: MMLU (Moral Scenarios) (Hendrycks et al., 2020), MMLU_PRO (Professional Law) (Wang et al., 2024), and AGIEval (LogicQA) (Zhong et al., 2023); Coding: MBPP (Austin et al., 2021); Reasoning: DROP (Dua et al., 2019) and ARC-Challenge (Clark et al., 2018); and Tool-using: API-BANK (Lv 3 problems) (Li et al., 2023). For further details, please refer to Table 15 in the Appendix.

### v3

# Efficiently Selecting Response Generation Strategy by Self-Aligned Perplexity for Fine-Tuning LLMs

@@ Abstract
Fine-tuning large language models (LLMs) typically relies on producing large sets of input-output pairs. Yet for a given question, there can be many valid outputs. In practice, these outputs are often derived by distilling knowledge from teacher models, and they can vary depending on the specific teacher model or prompting strategy employed. Recent findings show that how these training outputs are generated can significantly affect the performance of the fine-tuned model, raising an important question: how do we pick the best data generation method from among numerous possibilities? Rather than exhaustively training and evaluating on each candidate, this paper proposes a scalable approximate method that assesses a small subset of generated data to estimate its suitability for a specific target LLM. Our central idea is that effective outputs should be familiar to the target LLM. While previous work measures familiarity with perplexity, we find that perplexity might be suboptimal in characterizing “familiarity” through empirical analyses and practical observations. To address this, we introduce self-aligned perplexity, a novel metric capturing how closely candidate outputs adhere to the target LLM’s own style and reasoning patterns. In this way, we can identify the most effective generation strategy on a small sample, then apply it to produce the complete training set. We demonstrate that training on data generated by the chosen method yields significant improvements across diverse reasoning-focused benchmarks, particularly in cases where different candidate methods lead to highly divergent training outcomes. Our implementation is publicly available at https://github.com/XuanRen4470/SPPL.
Efficiently Selecting Response Generation Strategy by Self-Aligned Perplexity for Fine-Tuning LLMs
Xuan Ren††thanks: Equal contribution. University of Adelaide xuan.ren@adelaide.edu.au Qi Chen* University of Adelaide Lingqiao Liu1††thanks: Corresponding author. University of Adelaide lingqiao.liu@adelaide.edu.au
@@ 1 Introduction
Prior research has explored improving responses by adding details or rationales, such as structuring ground truth step by step (Hsieh et al., 2023; Ranaldi and Freitas, 2024), incorporating rationales, or enriching responses with additional information (Zhang et al., 2024; Kang et al., 2023; Li et al., 2022). However, recent studies (Ren et al., 2024; Yang et al., 2024) suggest that more details or converting responses to step by step style do not always improve performance and that alignment with the LLM’s linguistic style is crucial.
@@
In our experiment, we observe that no single response generation strategy works universally across tasks. Thus, we need to create a method to find out the most effective way to generate responses for each task, rather than a single method for all tasks.
@@
The concurrent works (Xu et al., 2024; Kim et al., 2024) attempt to predict the effectiveness of response generation methods by evaluating the entire training dataset. They generate full training datasets using each method and then estimate training effectiveness based on scores computed via algorithms or reward models. However, these approaches are computationally expensive and not scalable.
@@
However, can we predict the effectiveness of each data generation methods efficiently? We observe an interesting phenomenon that each response generation method produces responses with a consistent style, meaning that a small subset of generated examples can effectively represent the entire dataset. Based on this assumption, we propose an efficient ranking pipeline that evaluates a limited number of samples (e.g., 50) to assess the performance of each response generation strategy. This approach uses an alignment estimation function to assign scores to each strategy, enabling us to identify the best-performing method without the need for a full-dataset evaluation.
@@
Previous research (Ren et al., 2024) used perplexity to measure a model’s familiarity with candidate question-answer pairs, proposing that lower-perplexity responses for the same input tend to yield better training performance. However, we found several cases where perplexity-based filtering was ineffective. For instance, responses structured in a step-by-step or redundant style often exhibit low perplexity but do not necessarily improve training outcomes. While some candidate responses (e.g., step-by-step or redundant ones) may achieve low perplexity under the target LLM, the model itself rarely generates such responses when producing answers freely. This suggests that low perplexity does not always indicate alignment with the model’s inherent reasoning style. These findings suggest that perplexity can be "hacked" by response style. Thus, traditional perplexity alone is insufficient for selecting the best response generation strategy.
@@
To address this, we propose self-aligned perplexity, a refined metric for measuring a model’s familiarity with target responses. The key idea is that a model is most familiar with the data it generates itself. Leveraging this, we modify perplexity computation by incorporating model-generated responses as in-context examples. Specifically, we first have the model produce initial responses, which is then appended to the question as in-context examples. A prompt enforce the model to pay attention to these examples when computing perplexity, thereby altering the probability estimation of the candidate response. If the target response deviates significantly from the model’s own generated response—the one it is most familiar with—the model assigns it a lower probability, increasing its perplexity. Our experiments show that self-aligned perplexity outperforms traditional perplexity in selecting effective data generation strategies.
@@
In our experiments, we observe a strong correlation between the proposed indicator and the ranking of training dataset performance. Furthermore, we construct a pool of answer generation strategies and demonstrate that applying our selection criterion leads to significant performance gains compared to the baselines—especially in scenarios where different data-generation methods produce highly divergent outcomes
@@ 2 Related Works
There has been extensive research into what types of data yield the best training outcomes for large language models (LLMs). Previous studies have identified several factors that positively influence model training, such as adding complexity (Xu et al., 2023), adding details (Zhang et al., 2024; Kang et al., 2023; Li et al., 2022), adding diversity (Luo et al., 2023), augmenting ground-truth answers in a step-by-step manner (Hsieh et al., 2023; Ho et al., 2022; Magister et al., 2023; Fu et al., 2023; Ranaldi and Freitas, 2024), and ensuring correctness (Trinh et al., 2024; Ranaldi and Freitas, 2024). However, in practice, these metrics are challenging to measure for a given dataset, making it difficult to determine the quality of training data based on these criteria. Ren et al. (2024) found that familiarity, measured by perplexity, significantly impacts model training.
@@
Perplexity has been widely used for different purpose in prior research. Perplexity has been used to select prompts (Gonen et al., 2022), showing that prompts with lower perplexity generally lead to better performance in question-answering tasks. It has also been used for selecting pretraining datasets (De la Rosa et al., 2022), detecting AI-generated content (Xu and Sheng, 2024; Hu et al., 2020), and selecting instruction-tuning data from a database (Mekala et al., 2024). Li et al. (2024) modify the perplexity score and propose “IFD” (Instruction Following Difficulty), which is used to select a small pool of challenging data from the original dataset for efficient training. Researchers hypothesize that higher perplexity indicates more challenging data, which can be beneficial for teaching LLMs new knowledge. In addition, perplexity or confidence-based curricula have been explored for NMT (Kocmi and Bojar, 2017) and general text generation (Platanios et al., 2019), where harder (high-perplexity) data are introduced progressively to improve sample efficiency. Unlike these studies, our focus is on identifying the best strategy to generate target responses (y) for a given input (x), rather than selecting difficult (x, y) pairs for training language models.
Recent efforts have begun to ask which teacher model produces the most useful synthetic targets. Xu et al. (2024) introduce a Compatibility-Adjusted Reward (CAR) and judge its quality by the Spearman correlation between CAR scores and downstream accuracy on two instruction-following datasets, each evaluated with a single meta-prompt. Kim et al. (2024) study nine datasets spanning mathematics, coding, and general instructions; they correlate several corpus statistics with training gains and combine them with principal-component analysis to rank teacher models. Our study differs in four key respects. First, we estimate a strategy’s quality from only a small sample of its outputs, making synthetic data generation and evaluation far more affordable. Second, we target accuracy improvement, not just rank correlation. Third, we experiment on a much broader benchmark: 17 diverse tasks plus 6 Plan-Bench planning tasks. Fourth, we evaluate our method on datasets generated using diverse meta-prompts, explicitly accounting for prompt variability.
@@ 3.1 Problem Definition
Let be a set of candidate answer-generation strategies, where each strategy produces a response for an input . Our goal is to select the strategy that yields the most effective training data to fine-tune a target model . Since generating the full dataset via the API for every strategy is costly, we evaluate a small subset of size () to estimate how well each strategy’s outputs align with .
@@ 3.2 The Familiarity Hypothesis
The work in Ren et al. (2024) suggests that if the model is more “familiar” with a given response, then the model can learn better with the given response. In their work, perplexity, which is correlated to the likelihood of generating a response with the model, is used to measure this familiarity score. In our study, we argue that perplexity is sub-optimal to measure familiarity. We suggest that familiarity can be more precisely measured by this equation:
where is a semantic similarity measure between and a sample response drawn from the model . In plain language, it quantifies how similar a candidate response is to the range of answers that the model might generate. It is straightforward to demonstrate that when , i.e., when only if is exactly identical to , the function becomes equivalent to the likelihood , and hence equivalent to perplexity. using perplexity as a surrogate to measure familiarity fails to account for the variety of responses that may be semantically equivalent to a candidate response, thereby underestimating the familiarity. In practice, this results in assigning an excessively high perplexity to a good candidate response that the model might actually be familiar with, as evidenced by our empirical study in section 6.1.
@@ 3.3 Self-Aligned Perplexity
To evaluate the effectiveness of different response generation strategies, we first construct a small calibration set , consisting of the first examples (e.g., ) input questions. For each strategy, we generate one candidate response per question . Rather than directly measuring the likelihood of a response given a question, as in traditional perplexity, we compute its likelihood under a prompt augmented with in-context exemplars to better reflect stylistic alignment.
Specifically, for each input , we let the target model generate an initial prediction . Since these initial predictions may be incorrect, we evaluate them using the same task-specific script used at test time, and retain only those that are correct. These correct model-generated responses form a style pool .
To measure how well a candidate (generated by a specific strategy) aligns with the model’s own response style, we compute its perplexity under a prompt that includes two in-context exemplars randomly sampled from the style pool (see Appendix 6 for the prompt). Each exemplar comes from a different question than () to avoid leaking the answer.
This measurement corresponds to what we refer to as the self-aligned perplexity (SPPL) of . Compared to standard perplexity, which only conditions on the input question, SPPL incorporates stylistic guidance from the model’s own correct outputs. Lower SPPL indicates stronger alignment with the model’s preferred generation style.
For each response generation strategy , we compute the SPPL for all of its generated responses in , and average the scores to obtain a strategy-level value . We then rank all strategies by their average SPPL and select the one with the lowest score to generate the full training set.
@@ 4 Benchmark Construction
In this section, we show how we use different strategies (distinct prompts and teacher LLMs) in generating high-quality responses with different styles.
@@ 4.1 Target LLMs and APIs
We use Mistral-7B-instruct-V2 (Jiang et al., 2023), Llama3-instruct (Dubey and Abhinav Jauhri, 2024) and Qwen-2.5-7B-Instruct(Qwen et al., 2025) as the target language models . In this paper, we refer to Llama3-instruct, Mistral-7B-instruct-V2, and Qwen-2.5-7B-Instruct as Mistral7B, Llama3, and Qwen2.5, respectively. We use GPT-4o, MiniGPT-4o, and Claude 3.5 APIs as teacher models for response generation. Specifically, we use gpt-4o-mini-2024-07-18 and gpt-4o-2024-08-06 (OpenAI, 2023) from OpenAI, and claude-3-5-sonnet-20240620 (Anthropic, 2023) from Anthropic.
@@ 4.2 Datasets
We use English reasoning datasets referenced in the technical reports of LLaMA3 (Dubey and Abhinav Jauhri, 2024), Mistral (Jiang et al., 2023), and Qwen-2.5 (Qwen et al., 2025) (the three target models in our experiments). We select datasets with at least 650 examples that can be evaluated via accuracy. If a dataset lacks sufficient training data, we reconstruct it to contain at least 400 training, 50 validation, and 200 testing examples.
For datasets with subcategories (e.g., MATH, MMLU, MMLU_PRO, API_BANK, AGIEVAL), we choose the challenging subcategory (i.e., with the lowest reported accuracy). For example, we include moral scenarios from MMLU, Professional Law from MMLU_PRO, Level 3 problems from API_BANK, geometry from MATH, and LogicQA from AGIEVAL; we also incorporate the Algebra subcategory from MATH as in Ren et al. (2024).
@@
Main-experiment corpus.
In total, our datasets include: Mathematics: GSM8K (Cobbe et al., 2021), MATH (Algebra) and MATH (Geometry) (Hendrycks et al., 2021); Commonsense reasoning: PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), Hellaswag (Zellers et al., 2019), and ECQA (Aggarwal et al., 2021); Reading comprehension: BoolQ (Clark et al., 2019) and SQuAD (Rajpurkar et al., 2016); Aggregated benchmarks: MMLU (Moral Scenarios) (Hendrycks et al., 2020), MMLU_PRO (Professional Law) (Wang et al., 2024), and AGIEval (LogicQA) (Zhong et al., 2023); Coding: MBPP (Austin et al., 2021); Reasoning: DROP (Dua et al., 2019) and ARC-Challenge (Clark et al., 2018); and Tool-using: API-BANK (Lv 3 problems) (Li et al., 2023). More details are in Table 19 (Appendix).
PlanBench Extension.
We further evaluate the most challenging subtasks of PlanBench (Valmeekam et al., 2023)—those on which GPT-3 attains an accuracy below 20%. The subtasks comprise plan generation, plan optimization, plan verification, plan reuse, plan generalization, and replanning. Although we experimented with various prompt formats, Qwen consistently failed to solve any execution problems. Since our method relies on generating correct responses for use as in-context examples, we exclude the execution task from our evaluation. The remaining six categories remain sufficiently challenging and are not part of our main training benchmarks. We use them solely to analyze performance variance when models are trained on tasks that are very challenging. How do different response generation strategies affect performance variance under such conditions?